OpenStudy (anonymous):

A line passes through the points (-10,-4) and (-1,2). What is the slope of the line that passes through the two points?

OpenStudy (anonymous):

slope is given by (y2 - y1)/(x2 - x1) where (x1,y1) and (x2,y2) are the points :)

OpenStudy (anonymous):

I don't understand....

OpenStudy (anonymous):

Can you help me understand it a little better?

OpenStudy (aravindg):

For example take points (3.4) ,(5,6) Slope=y difference/x difference =(6-4)/(5-3) =2/2=1

OpenStudy (aravindg):

Apply the same concept here.